OpenStudy (anonymous):

The sum of three consecutive even integers is six more than two times the smallest integer. Find all the integers.

OpenStudy (anonymous):

Let "x" be the smallest integer this would mean the next even integer can be described by "x+2" what would the next integer be represented as?

OpenStudy (anonymous):

write even integers as 2k

OpenStudy (anonymous):

three consecutive even numbers would be 2k, 2(k+1), 2(k+2)

OpenStudy (anonymous):

see if you can use that to try and write an equation from the statement you are given

OpenStudy (anonymous):

and solve for k

OpenStudy (cruffo):

I would suggest something more along the lines of the strategy that @completeidiot was using.

OpenStudy (cruffo):

For example, three consec. even entegers: 2, 4, 6 If you let x be the smallest of the three integers: x = 2 then you can get to the next number by adding 2 to x : x+2 = 2+2 = 4 and then the last number is just x+4: x+4 = 2+4 = 6

OpenStudy (cruffo):

Of course the number 2,4,6 is not the actual answer to your question, but small cases like this may help you understand the problem.

OpenStudy (cruffo):

Question: The sum of three consecutive even integers is six more than two times the smallest integer. Translate: x + (x+2) + (x+4) = 2x+6

OpenStudy (anonymous):

@cruffo no, specifying a number as 'x' doesn't guarantee that it will be even.

OpenStudy (anonymous):

that's why you use 2k, 2(k+1), etc.

OpenStudy (cruffo):

@Algebraic!, I understand what you are saying by "2k". Your method will also work. However, it is not actually necessary for these types of number problems.

OpenStudy (cruffo):

In your method, you get 2k + 2(k+1) + 2(k+2) = 2(2k) + 6 and you can divide out all the 2's

OpenStudy (cruffo):

then you get k + k+1 + k+2 = 2k + 3

OpenStudy (cruffo):

However solving for k won't give you the smallest integer. after you solve for k, you'll need to go back and multiply k by 2 to get the final answer.

OpenStudy (cruffo):

Actually, I don't think you actually get an answer by doing that: 3k + 3 = 2k+3 k=0???

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

ask cruffo then. the problem is already done, however.